OpenStudy (anonymous):

Solve x2 – 5x – 1 = 0 using the Quadratic Formula. Find the exact solutions. Then approximate any radical solutions. Round to the nearest hundredth.

Directrix (directrix):

To get started see attached Quadratic Formula

OpenStudy (jhannybean):

what is your a, b, and c?

OpenStudy (anonymous):

i am not sure

OpenStudy (danjs):

The general form of your equation is Ax^2 + Bx + C = 0 A=1 B=-5 C=-1 Use those and put them into the equation that was linked... what do you get so far?

OpenStudy (danjs):

I plugged in the numbers for you. Just have to simplify now. \[x = \frac{ -B \pm \sqrt{B^2 - 4*A*C} }{ 2*A } = \frac{ 5 \pm \sqrt{(-5)^2 -4 * 1 *(-1)} }{ 2 * 1 }\]

OpenStudy (anonymous):

how do i get the correct symbols?

OpenStudy (jhannybean):

What do you mean exactly?

OpenStudy (danjs):

just add and multiply numbers

OpenStudy (anonymous):

never mind i thought i need all the square and plus or minus symbols but i get it now XD

OpenStudy (danjs):

x=\[\frac{ 5 \pm \sqrt{29} }{ 2 }\] right? or did i mess up

OpenStudy (jhannybean):

inside the square root: 25 + 4

OpenStudy (anonymous):

yup thats it @DanJS thank you!

OpenStudy (danjs):

you're welcome, anytime

OpenStudy (anonymous):

i think im doing it wrong i keep getting 6.35

OpenStudy (anonymous):

oh on my answers it says divide by 2 not 4?

OpenStudy (danjs):

oh right, that was a typo on my part

OpenStudy (anonymous):

but im still not getting the right answer

OpenStudy (danjs):

\[\frac{ 5 - \sqrt{29} }{ 2 } \approx -0.193\]

OpenStudy (danjs):

\[\frac{ 5 + \sqrt{29} }{ 2 } \approx 5.193\]

OpenStudy (anonymous):

would i put the + or + first?

OpenStudy (danjs):

they are 2 separate solutions, one is plus the other is minus, like my above 2 posts

OpenStudy (danjs):

the graph crosses the x axis at 2 values, those satisfy the equation in the question

OpenStudy (anonymous):

i found out the one answer had a -5 haha tricky tricky tricky thank you!

OpenStudy (danjs):

welcome, anytime